A vapor stream that is 65 mole% styrene and 35 mol
Answer #1

Use Raoult's law to solve this problem,
You have to find XS , Xt and T for this problem
Ys =0.65 (vapor mole fraction of styrene) ; YT = 0.35 (vapoor mole fraction of toluene)
Xs = liquid mole fraction of styrene ; XT = liquid mole fraction of toluene
T = temperature for the system
So, you need three equations to solve for 3 unknowns ( XS , Xt and T)
YsPtotal = XsPs
0.65 * 150 mm Hg =  XsPs
97.5 / Ps  = Xs --------------- (1)
(1- Ys) Ptotal = (1 - Xs)PT   
0.35 * 150 mmHg = (1 - Xs)PT   
52.5 = (1 - Xs)PT ( 1-Xs = XT)
52.5 / PT = XT --------------------- (2)
XT + Xs =1
(97.5 / Ps) + (52.5 / PT ) = 1 --------------- (3)
( Ps in mm Hg ,T in oC)
log Ps = 4.21948 - (1525.059) / T - 56.379    ----------- (4)
log PT = 4.07827 - (1343.943) / T - 53.773 (For toluene) --------------- (5)
solve equations 3, 4 & 5 to get Ps ,PT &T
Now from Ps and PT find Xs and XT from equations 1 & 2